# Foveated Retinotopy Improves Classification and Localization in CNNs

### v3

@@ Results
We retrain the pre-trained ResNet [33] networks on different variants of the ImageNet-1K dataset [32] (see Methods “Datasets and Dataset Transformations”), either using a simple circular mask (hereafter called the Cartesian dataset) or using log-polar transformed images (hereafter called the “retinotopic” dataset). Two variants of the training sets are also considered. In a first case (called the “regular” case), the mask or the log-polar transformed images are applied to the “regular” images. In a second case (called the “focus” case), the mask or the log-polar transformation is focused on the center of the bounding boxes surrounding the objects of interest provided with the data set. The pre-trained networks are first re-trained on the “regular” version of the dataset, generating a first generation of networks, and then a second generation is trained by fine-tuning these networks on the “focus” dataset. We then compare the classification accuracy of the original and the different re-trained networks.
The baseline classification accuracy of the standard pre-trained ResNet network, which we will call the “raw” network (as it has no retinotopic transformation nor training), averaged on the Cartesian test dataset is $81.7\%$. In comparison, our re-trained networks show respectively accuracies of $78.5\%$ and $74.3\%$ on the “regular” Cartesian and retinotopic datasets, and $82.5\%$ and $77.4\%$ on the “focus” Cartesian and retinotopic datasets. As we can see, when fine-tuning the networks on the “focus” dataset, despite reducing the image resolution by cropping the image, the accuracy is improved by $4.0\%$ and $3.1\%$ respectively, demonstrating that this cropping tends to suppress distractors in the periphery and enhances average accuracy.
Furthermore, our networks re-trained on the log-polar transformed images have a slightly lower categorization accuracy than those re-trained on Cartesian images (see Figure 2-A & Figure 3-A). This result was expected because the log-polar transformation discards fine-grained information in the periphery without increasing central resolution. In fact, this rather limited loss of accuracy is quite remarkable for such a massive loss of information. The reason could be a general “photographer’s bias” that tends to place the main object in the central region of the image.
@@
To what extent can these re-trained networks be relied upon in practical applications? Indeed, a persistent challenge for deep neural networks is their lack of reliability in critical situations. For example, it is well documented that a minor alteration to the inputs of these networks, if well-designed, may result in a significant decline in classification accuracy. In particular, a classical robustness test applied to deep networks, called an “adversarial attack,” consists of perturbing independently the pixels in each image in order to maximize the error rate in the test phase (see the Methods “Attacking classical CNNs with a geometrical rotation”). However, these modifications are typically subtle and perceptually resemble identically distributed independent noise, rendering them unlikely to happen in natural conditions. In contrast, biological agents interacting with their environment undergo significant visual perturbations, including large scene pans and tilts due to head and body movements.
@@
As real-world objects appear in different orientations, we assess the resilience of our re-trained networks to geometric alterations of their inputs e.g. rotations and zooms. In comparison to modifying individual pixels, a rotation or a zoom represents a coherent, whole-image transformation controlled by a single parameter, namely the rotation angle or scaling factor. We thus choose to investigate an attack that maximizes the classification drop for each image individually, and then compute the corresponding accuracy averaged over a sample of $50,000$ images from the ImageNet validation dataset.
@@
Fig 2: (A) For all the networks, we plot the accuracy averaged over the dataset without rotation (in light color), or for each image rotated at the angle $\bar{\theta}$ with the worst loss (rotation-based attack, in full color). No shading : regular dataset. Diagonal shading : focus dataset. Gray dashed lines: accuracy of the “raw” network (without any transformation nor training). (B) The average accuracy is shown for both Cartesian or retinotopic re-trained networks, and the “raw” network, with different image rotations. The rotation is applied around the central fixation point with an angle ranging from $-180^{\circ}$ to $+180^{\circ}$ (in steps of $15^{\circ}$).
@@
Our experiments show that while the original ResNet $101$ achieves a nominal baseline accuracy of $81.7\%$ on unperturbed images, a rotation attack significantly reduces the performance of the model. Applying the maximally deceptive rotation to each image reduces the average accuracy to $46.1\%$ (see gray lines in Figure 2-A). This is also true for our Cartesian network, with accuracy dropping to $39.3\%$ for the Cartesian network re-trained on “regular” images, and even to $37.4\%$ for the network fine-tuned on the “focus” dataset. In contrast, the retinotopic networks show a lower sensitivity to rotation attacks, with an accuracy reduced to $47.0\%$ when using the “regular” images, and even to $55.4\%$ when using the “focus” dataset.
@@
This difference between the two types of networks is even more manifest in Figure 2-B, where we computed the average accuracy over the test dataset for each rotation. As expected, the Cartesian networks show a monotonic decrease in accuracy with respect to the angle of rotation, and a symmetry with respect to horizontal flips. Strikingly, this effect is nearly absent in our retinotopic networks (see Figure 2-B), which show a flat (invariant) accuracy over the whole range of rotation angles. This marked difference can be interpreted as a consequence of the horizontal translation invariance found in classical CNNs. When applied to the retinotopic input space, this invariance transforms seamlessly into rotation invariance in visual space [22] (see [42] for a proof).
@@
Fig 3: (A) For all networks, we plot the accuracy averaged over the unzoomed dataset (in light color), or for each image zoomed out to the scale with the worst loss (zoom out attack, in full color). The gray dashed lines represent the accuracy of the “raw” network (without re-training). (B) Average accuracy over a sample of $50,000$ images from the ImageNetvalidation dataset, shown for both re-trained and pre-trained networks with different input image zooms. The zoom is applied at the central fixation point with a zoom factor ranging from ×$\times 10$ to ×$\times 0.1$.
@@
Analogous to rotation, zooming in and out is equivalent to a translation in log-polar space, and this property is expected to induce a similar invariance in the retinotopic networks. Similar experiments were therefore performed to test the effect of a zoom (see Figure 3), with a zoom ranging from ×$\times 10$ to ×$\times 0.1$, divided into “zoom-in” (×$\times 10$ to ×$\times 1$ range) and “zoom-out” (×$\times 1$ to ×$\times 0.1$ range, in steps of $0.1$). In Figure 3-A, we applied the zoom attack only for the zoom-out case, i.e. ratios between ×$\times 1$ and ×$\times 0.1$, because zoom-in attack scenarios always lead to the maximum zoom-in, where the accuracy approaches zero. All Cartesian networks response (i.e. “raw”, “regular” or “focus”) becomes close chance-level to when submitted to a zoom-out attack. Only the retinotopic networks keep a discriminative capacity in this case ($51.0\%$ on “regular” images and $53.6\%$ on “focus” images), illustrating the importance of foveal information in categorization, especially when the peripheral information is scarce or deceptive.
@@
Figure 3-B shows the average accuracy over the entire zoom range, providing a more comprehensive view of the effect of zooming in and out on our different networks. Contrary to the previous case, a clear asymmetry of the effect is observed in both cases. Specifically, the different Cartesian networks tested here (the original and the re-trained) show an approximate symmetry with respect to the logartihm of zoom-in/zoom-out factors. In contrast, the retinotopic frame shows resilience to changes in zoom level when zooming out but shows a significant loss of accuracy, even stronger than for the Cartesian networks, when zooming in. Finally, both retinotopic and Cartesian networks show a sensitivity to zoom-in transformations, with their mean accuracies dropping close to chance levels at a ×$\times 10$ zoom. However, this observation must be treated with caution, as it is likely due to the loss of visual information and blurring that occurs when extreme zoom levels are applied to these relatively low-resolution images. Interestingly, the Cartesian networks achieve slightly higher accuracies here, suggesting a potential advantage for this architecture under zoom-in changes.
@@
In contrast to the case of rotations, our fine-tuning on the “focus” dataset has a mixed impact here, as evidenced by the zoom optimal point. With the Cartesian network, either with or without re-training, the zoom corresponding to the maximum accuracy appears to be ×$\times 2$, undergoing a shift to ×$\times 1$ when fine-tuned on the “focus” dataset. The optimal zoom for the retinotopic frame remains constant at ×$\times 1$ on either “regular” or “focus” data. From an ecological perspective, zoom-out invariance is probably one of the most valuable features in natural vision. It enables the detection of predators or prey at any scale, even in complex and crowded environments where survival-critical information may appear at different sizes. This ability allows for a more accurate interpretation of the visual scene but comes with a trade-off: the need to position objects of interest at the center of the retina.
@@
Fig 4: (A) For all networks, we plot the accuracy averaged over the validation dataset (in light color), or for each image translated (rolled) to the position with the worst loss (translation attack, in full color). The gray dashed lines represent the accuracy of the “raw” network (without re-training). (B) Average accuracy over a sample of $50,000$ images from the ImageNet [32] validation dataset, shown for both re-trained and pre-trained networks with different input image translation roll. The translation is applied from the central fixation point and defined a linear grid of $11\times 1$ 11 point of fixations
@@
Finally, based on these observations and the fact that translations in Cartesian space induce a significant, nonlinear transformation in retinotopic space (see Figure 1-A&B), we investigate the effect of translations in Cartesian space on retinotopic networks. We thus apply a roll function to the input image to place different fixation points in the center. The fixation points are linearly distributed on an $11\times 1$ 11 grid. We then plot the mean accuracy when systematically selecting the worst position (based on the minimum loss as in the rotation Figure 2-A or zoom-out attack Figure 3-A) for the target label (see Figure 4-A), or the mean accuracy of the networks as a function of the position repositioned in the center (see Figure 4-B).
@@
Visual object localization : likelihood Heatmaps
@@
To quantify the contribution of this translation sensitivity, we consider here a new task, i.e. the visual search task, in which a visual object (of which the label is known in advance) needs to be localized over the entire image. Is is known that such task allows for the quick retrieval of an image [43]. This protocol allows us to compare different visual shifts, each one corresponding to a potential fixation point, and to generate a map of the expected (or actual) accuracy, for each network, as a function of the fixation point (see Methods “Visual object localization : Protocol”). In this protocol, a set of $11\times 1$ 11 fixation points is defined (with the coordinate $(5,5)$ being the center of the image), and at each coordinate of the grid, a likelihood value is computed for the label of interest (see figure 13). In practice, a likelihood is given at each location from a softmax calculation over the different labels, as in the classical ResNet classifier, providing a value between 00 and $1$ for the label of interest. This projects the network output onto a $11\times 1$ 11 Bernoulli probability space corresponding to the likelihood of detecting the given label at each position, finally providing a “heat map” on our $11\times 1$ 11 grid in a way that is compatible with other localization protocols. We tested our likelihood protocol on the “regular” validation dataset (see Fig 5 for some examples).
@@
Fig 5: Likelihood maps computed on typical image using $11\times 1$ 11 fixation points with the “raw” network, that is the original ResNet classifier with no re-training (second column), the network re-trained on images with circular mask on the “focus” dataset (third column), and the one re-trained on log-polar tranformed inputs (last column). The map displays the likelihood for the label of interest in the image.
@@
To further quantify the difference between the Cartesian and retinotopic cases, we considered the localization information provided by the bounding boxes. Indeed, the bounding boxes separate the image into two regions, a region where the object of interest is present and a region where it is absent. Thus, from each likelihood map in the validation set, we computed the mean likelihood for the label under study within the bounding boxes (ground truth from the dataset) and the mean likelihood outside these boxes. Given the higher classifications rates of the networks when fine-tuned on bounding boxes only the “focus” networks are used for the remainder of the study.
The results are shown in Figure 6-A. The figure shows a mean likelihood value obtained inside and outside the bounding boxes for the original ResNet network, our re-trained Cartesian network, and the re-trained retinotopic network. Both networks show significantly higher likelihoods when the fixation point is inside the bounding box than when it is outside the bounding box. This reflects a higher confidence in the label response. At first glance, the average likelihoods in the three cases seem quite comparable, although they are slightly higher in the re-trained Cartesian case. This confirms the presence of a residual translation sensitivity in the Cartesian CNNs, even if this sensitivity is more pronounced in the retinotopic case.
@@
Fig 6: We tested the visual search protocol on the ImageNet validation dataset ($50,000$ images). (A) The mean likelihood across the point of fixation inside the bounding box (“In”) or the point of fixation outside the bounding box (“Out”) and the corresponding ratio of activation. (B) The intersection over union as a function of a threshold applied on the likelihood map.
@@
In Figure 6-B we use an additional method: We consider the Intersection over Union (IoU) to evaluate the agreement between the bounding boxes and the activation maps (see Methods “Localization tools and evaluation”). The peak threshold is set close to 00 for all versions of the networks. We find that the Cartesian network consistently shows higher IoU values compared to the retinotopic network. Furthermore, the networks fine-tuned on the “focus” datasets introduce a noticeable change: both networks show a slower decay rate, leading to a stronger correlation with the ground truth.
Visual object localization : Mean likelihood maps
A complementary method of comparison is used in Figure 7-A, which shows the mean likelihood of the label of interest as a function of the distance from the most salient position. The maps were re-centered to place the viewpoint with the highest likelihood at the center of the fixation map. During the re-centering process, the spots outside the image boundary were assigned a “Nan” value to facilitate boundary management. Accordingly, the mean maps were generated using the “nanmean()” function, and the mean activation is displayed.
Fig 7: Top row : The following maps illustrate the mean likelihood of the investigated networks when tested on the validation ImageNet dataset ($50000$ images). First row: recentered maps, averaged over the validation dataset. From left to right: “raw” network (no re-training), Cartesian network on the “focus” dataset , and Retinotopic network on the “focus” dataset. Middle row :  Difference maps. From left to right: Cartesian minus “raw”, Retinotopic minus “raw”, Cartesian minus retinotopic. Bottom row :  Log-odd ratio maps, same combinations.
Our recentered likelihood maps are shown on fig.7. All three maps show a similar 2D bell-shaped activation, indicating a clear object positioning capability for all networks, retinotopic or not, when applying our “visual search” protocol (i.e. without rolling the image borders). The activation level is higher at the center (“near” the object of interest), and lower at the periphery. The likelihood values are different though, being higher on average for the Cartesian network, lower on average for the “raw” network, and more contrasted for the retinotopic one.
To further quantify this contrast difference, we calculate the map difference on the second row of fig. 7, along with the log-odd ratio in the third row of Figure 7, considering three different cases, i.e. Cartesian vs. “raw”, retinotopic vs. “raw”, and Cartesian vs. retinotopic.
In the second row, the difference between the Cartesian networks and the “raw” network shows only positive values, while the difference between the retinotopic networks and the “raw” network has both negative and positive values. The difference between the retinotopic and Cartesian networks has all values negative, indicating on average a higher likelihood level in the Cartesian case. The area around the center remains close to zero though, reflecting a sharper slope towards the peripheral region in the retinotopic case.
The log-odd ratio maps (third row) look quite similar to the difference maps, except in the center where the differences are more manifest. Of particular interest here is the comparison of the first and second column, i.e. Cartesian vs. “raw” (first column), and Retinotopic vs. “raw” (second column). The radius of the central spot appears clearly different in the two cases, with a large central spot in the Cartesian case, reflecting a loosy spatial discrimination improvement, and a more reduced one in the retinotopic case, reflecting a more acute and object-centric spatial discrimination improvement. Moreover, only the retinotopic network (second column) has significant peripheral depletion in comparison with the “raw” network. These observations collectively support the idea that the retinotopic network generates a heat map with enhanced contrast around the area of interest. In addition though, a focal Cartesian network also manifests object localization capabilities, with high likelihood values around the object of interest, though being less spatially specific.
Visual object localization : Heatmap accuracy
In this section, we consider both the localization and categorization capabilities of our networks. For a given label, known in advance, what is the actual accuracy obtained when placing the eye at the position of highest likelihood, i.e. which improvement can be expected, with respect to the baseline, when placing the eye (or camera) at its optimal position? In that case, the accuracy of all networks increases to about $96\%$ (“raw” $97.7\%$, Cartesian $96.2\%$, retinotopic $95.8\%$, see Figure 8). This massive improvement generally reflects the critical role of object centering in classification, even in the Cartesian case, an issue rarely evoked in mainstream computer vision. Moreover, in addition to the highest contrast observed in the likelihood maps, the improvement is also stronger in the retinotopic case, with $18.7\%$ improvement in the “raw” case, $18.9\%$ in the Cartesian case and $22.9\%$ improvement in the retinotopic case, see Figure 8). Also, because the networks are fine-tuned on the boxes, their pre-saccadic acuity is a discrete test, since they are presented with the full off-center image.

### v4

@@ Results
We retrain pre-trained ResNet (He et al., 2015) networks on different variants of the ImageNet-1K dataset (Russakovsky et al., 2015) (see Methods “Datasets and Dataset Transformations”), either using a simple circular mask applied on the raw images (hereafter called the Cartesian dataset) or using log-polar transformed images (hereafter called the “retinotopic” dataset). Two variants of the training sets are also considered. In a first case (called the “regular” case), the mask or the log-polar transformed images are applied to the “regular” images. In a second case (called the “focus” case), the mask or the log-polar transformation is focused on the center of the bounding boxes surrounding the objects of interest and which are provided with the data set. The pre-trained networks are first re-trained on the “regular” version of the dataset, generating a first generation of networks, and then a second generation is trained by fine-tuning these networks on the “focus” dataset (see Figure 11). We then compare the classification accuracy of the original and the different re-trained networks. In this section we choose to express the accuracy as the percentage of correct categorization on the corresponding validation dataset.
The baseline classification accuracy of the standard pre-trained ResNet network, which we will call the “raw” network (as it has no mask, retinotopic transformation nor training), averaged on the Cartesian test dataset is $81.7\%$. In comparison, our re-trained networks show respectively accuracies of $78.5\%$ and $74.3\%$ on the “regular” Cartesian and retinotopic datasets, and $82.5\%$ and $77.4\%$ on the “focus” Cartesian and retinotopic datasets. As we can observe, when fine-tuning the networks on the “focus” dataset, despite reducing the image resolution by cropping the image, the accuracy is improved by $4.0\%$ and $3.1\%$ respectively. This demonstrates that this cropping tends to suppress distractors in the periphery and enhances average accuracy.
Furthermore, our networks re-trained on the log-polar transformed images have a slightly lower categorization accuracy than those re-trained on Cartesian images (see Figure 2-A & Figure 3-A). This result was expected because the log-polar transformation discards fine-grained information in the periphery without increasing central resolution. In fact, this rather limited loss of accuracy is quite remarkable for such a massive loss of information. One reason could be a general “photographer’s bias” for the “regular” images that tends to place the main object in the central region of the image. It is confirmed by the increase of accuracy observed when centering the point of fixation in the “focus” dataset.
@@
To what extent can these re-trained networks be relied upon in practical applications? Indeed, a persistent challenge for deep neural networks is their lack of reliability in adversarial situations. For example, it is well documented that a minor alteration to the inputs of these networks, if well-designed, may result in a significant decline in classification accuracy. In particular, a classical robustness test applied to deep networks, called an “adversarial attack,” consists of perturbing independently the pixels in each image in order to maximize the error rate in the test phase (see the Methods “Attacking classical CNNs with a geometrical rotation”). However, these modifications are typically subtle and perceptually resemble identically distributed independent noise, rendering them unlikely to happen in natural conditions. In contrast, biological agents interacting with their environment undergo significant visual perturbations, including large scene pans and tilts due to head and body movements.
@@
As real-world objects appear in different orientations, we assess the resilience of our re-trained networks to geometric alterations of their inputs e.g. rotations and zooms. In comparison to modifying individual pixels, a rotation or a zoom represents a coherent, whole-image transformation controlled by a single parameter, namely the rotation angle or scaling factor. We thus choose to investigate for each network a simple attack scenario that maximizes the loss for different rotation angles for each image individually and then evaluates the classification result for this “worst” angle. Then, we compute the accuracy averaged over a sample of $50,000$ images from the ImageNet validation dataset.
@@
Figure 2: (A) For all the networks, we plot the accuracy averaged over the dataset without rotation (in light color), or for each image rotated at the angle $\bar{\theta}$ with the worst loss (rotation-based attack, in full color). No shading : regular dataset. Diagonal shading : focus dataset. Gray dashed lines: accuracy of the “raw” network (without any transformation nor training). (B) The average accuracy is shown for both Cartesian or retinotopic re-trained networks, and the “raw” network, with different image rotations. The rotation is applied around the central fixation point with an angle ranging from $180^{\circ}$ to $180^{\circ}$ (in steps of $15^{\circ}$).
@@
Our experiments show that while the original ResNet $101$ achieves a nominal baseline accuracy of $81.7\%$ on unperturbed images, a rotation attack significantly reduces the performance of the model. Applying the maximally deceptive rotation to each image reduces the average accuracy to $46.1\%$ (see gray lines in Figure 2-A). This is also true for our Cartesian network, with accuracy dropping to $39.3\%$ for the Cartesian network re-trained on “regular” images, and even to $37.4\%$ for the network fine-tuned on the “focus” dataset. In contrast, the retinotopic networks show a lower sensitivity to rotation attacks, with an accuracy reduced to $47.0\%$ when using the “regular” images, and only to $55.4\%$ when using the “focus” dataset.
@@
This difference between the two types of networks is even more manifest in Figure 2-B, where we compute the average accuracy over the test dataset for each single rotation angle. We first observe that the Cartesian networks show a decrease in accuracy with respect to the angle of rotation (with a symmetry with respect to horizontal flips), with a significant drop around $160^{\circ}$, “raw” and “focus” accuracies were degraded to $66.8\%$ and $68.1\%$, i.e. a drop of $16.1\%$ and $13.4\%$ ($64.3\%$ in “regular”, i.e. a drop of $15.8\%$). Strikingly, this effect is nearly absent in our retinotopic networks (see Figure 2-B), which show a flat (invariant) accuracy over the whole range of rotation angles, with a minimal degraded accuracy at $76.4\%$, i.e. a drop of $0.8\%$ ($71.6\%$ in “regular”, i.e. a drop of $2.9\%$). This marked difference can be interpreted as a consequence of the horizontal translation invariance found in classical CNNs. When applied to the retinotopic input space, this invariance transforms seamlessly into rotation invariance in visual space (Araujo and Dias, 1997) (see (Hao et al., 2021) for a proof). Note that for all networks, no maximum accuracy exceeds the value obtained without rotation (i.e. $0^{\circ}$ rotation angle) by more than $0.01\%$.
@@
Figure 3: (A) For all networks, we plot the accuracy averaged over the unzoomed dataset (in light color), or for each image zoomed out to the scale with the worst loss (zoom out attack, in full color). The gray dashed lines represent the accuracy of the “raw” network (without re-training). (B) Average accuracy over a sample of $50,000$ images from the ImageNet validation dataset, shown for both re-trained and pre-trained networks with different input image zooms. The zoom is applied at the central fixation point with a zoom factor geometrically-spaced from $\times 10$ to $\times 0.1$.
@@
Analogous to rotation, zooming in and out is equivalent to a translation in log-polar space, and this property is expected to induce a similar invariance in the retinotopic networks. Similar experiments were therefore performed to test the effect of a zoom (see Figure 3), with a zoom ranging from $\times 10$ to $\times 0.1$, divided into “zoom-in” ($\times 10$ to $\times 1$ range) and “zoom-out” ($\times 1$ to $\times 0.1$ range). In Figure 3-A, we applied the zoom attack only for the zoom-out case, i.e. ratios between $\times 1$ and $\times 0.1$, because zoom-in attack scenarios always lead to the maximum zoom-in, where the accuracy approaches zero. The response of all Cartesian networks (i.e. “raw”, “regular” or “focus”) reach chance-level performance when submitted to a zoom-out attack. Only the retinotopic networks keep a discriminative capacity in this case ($51.0\%$ on “regular” images and $53.6\%$ on “focus” images), illustrating the importance of foveal information in categorization, especially when the peripheral information is scarce or deceptive.
@@
Figure 3-B illustrates the average accuracy across the entire zoom range, offering a comprehensive perspective on the effects of zooming in and out across our different networks. Unlike the previous case, a pronounced asymmetry emerges in both scenarios. Specifically, the various Cartesian networks tested (both the original and re-trained versions) exhibit an approximate symmetry relative to the logarithmic scale of zoom-in and zoom-out factors. In contrast, the retinotopic network displays an asymmetric response, characterized by a slightly declining plateau in the zoom-out direction and a more pronounced accuracy loss in the zoom-in direction compared to the Cartesian networks. When zooming in, both the retinotopic and Cartesian networks exhibit heightened sensitivity, with mean accuracy dropping near chance levels at a $\times 10$ zoom. Notably, the Cartesian networks achieve marginally higher accuracies in this scenario, suggesting a potential architectural advantage under extreme zoom-in conditions. However, this finding should be interpreted cautiously, as it likely stems from the loss of visual detail and the blurring effect inherent to applying extreme zoom levels on low-resolution original images.
@@
In contrast to the case of rotations, our fine-tuning on the “focus” dataset has a nuanced effect, as evidenced by its impact on the zoom optimal point. For the Cartesian network, the zoom corresponding to maximum accuracy is around $\times 2$ in both its pre-trained and re-trained states. However, fine-tuning on the “focus” dataset shifts this optimal zoom to $\times 1$. In comparison, the optimal zoom for the retinotopic network remains consistently at $\times 1$, regardless of evaluation on “regular” or “focus” data. From an ecological perspective, zoom-out invariance is likely one of the most advantageous features in natural vision. This capability facilitates the detection of predators or prey at a distance, even in complex and cluttered environments where survival-critical information may appear at varying sizes. While this feature enhances the interpretation of visual scenes, it comes with a trade-off: the necessity to position objects of interest centrally on the retina.
@@
Figure 4: (A) For all networks, we plot the accuracy averaged over the validation dataset (in light color), or for each image translated (rolled) to the position with the worst loss (translation attack, in full color). The gray dashed lines represent the accuracy of the “raw” network (without re-training). (B) Average accuracy over a sample of $50,000$ images from the ImageNet validation dataset, shown for both re-trained and pre-trained networks with different input image translation rolls. The translation is applied from the central fixation point and defines a linear grid of $11\times 11$ points of fixation distributed regularly on the image.
@@
Finally, based on these observations and the fact that translations in Cartesian space induce a significant, nonlinear transformation in retinotopic space (see Figure 1-A&B), we investigate the effect of translations in Cartesian space on retinotopic networks. We thus investigate the effect of a rigid full-field translation by applying a roll function to the input image and place the fixation point at different positions in the image. Specifically, the fixation points are linearly distributed on an $11\times 11$ grid. We then plot the mean accuracy when systematically selecting the worst position (based on the minimum loss as in the scenarios for a rotation Figure 2-A or zoom-out attack Figure 3-A) for the target label (see Figure 4-A), or the mean accuracy of the networks as a function of the position repositioned in the center (see Figure 4-B).
@@
Visual object localization : Likelihood maps
Protocol
@@
To quantify the contribution of this sensitivity to translation, we consider here a new task, i.e. the visual search task, in which a visual object (of which the label is known in advance) needs to be localized over the entire image. Is is for instance known that such task allows for the quick retrieval of an image label (Potter, 1975). We design a protocol for each network to allow us to compare different visual shifts, each one corresponding to a potential fixation point, and to generate a map of the expected (or actual) accuracy as a function of the fixation point (see Methods “Visual object localization : Protocol”). In this protocol, a set of $11\times 11$ fixation points is defined (with the coordinate $(5,5)$ being the center of the image), and at each coordinate of the grid, a likelihood value is computed for the label of interest (see figure 13). In practice, a likelihood is given at each location from a softmax calculation over the different labels, as in the classical ResNet classifier, providing a value between 0 and $1$ for the label of interest. This projects the network output onto a $11\times 11$ Bernoulli probability space corresponding to the likelihood of detecting the given label at each position, finally providing a “heat map” on our $11\times 11$ grid in a way that is compatible with other localization protocols. We tested our likelihood protocol on the “regular” validation dataset (see Fig 5 for some examples).
@@
Figure 5: Likelihood maps computed on some prototypical images using $11\times 11$ fixation points with the “raw” network, that is the original ResNet classifier with no re-training (second column), the network re-trained on images with circular mask on the “focus” dataset (third column), and the one re-trained on log-polar tranformed inputs (last column). The map displays the likelihood for the label of interest in the image.
Mean likelihood maps
In Figure 6, we calculated the likelihood maps for all the images of the validation set, and re-centered them to place the viewpoint with the highest likelihood at the center of the grid. During the re-centering process, the spots outside the grid were assigned a Nan value to facilitate boundary management, and the average “center maps” were generated using the nanmean() function, providing the likelihood profile of the label of interest as a function of the distance from the most salient position. Our recentered likelihood maps are shown on Figure 6. All three maps show a similar 2D bell-shaped activation, indicating a clear object positioning capability for all networks, retinotopic or not, when applying our “visual search” protocol (i.e. without rolling the image borders). The activation level is higher at the center (“near” the object of interest), and lower at the periphery. The likelihood values are different though, being higher on average for the Cartesian network, lower on average for the “raw” network, and more contrasted for the retinotopic network. To further quantify this contrast difference, we calculate the map difference on the second row of Figure 6, along with the log-odd ratio in the third row of Figure 6, considering one-to-one comparisons, i.e. Cartesian vs. “raw”, retinotopic vs. “raw”, and Cartesian vs. retinotopic.
Figure 6: Mean likelihood on ImageNet’s validation dataset ($50000$ images). From left to right: “raw” network (no re-training), Cartesian network retrained on the “focus” dataset, and Retinotopic network retrained on the “focus” dataset. First row: recentered maps, averaged over the validation dataset. Middle row: Difference maps. Bottom row: Log-odd ratio maps.
Let’s analyze Figure 6. The second row shows the difference maps. The Cartesian minus “raw” map shows only positive values, while the retinotopic minus “raw” network has both positive and negative values. The retinotopic minus Cartesian map has only negative values, indicating on average a higher likelihood level in the Cartesian case. The area around the center remains close to zero though, reflecting a sharper slope towards the peripheral region in the retinotopic case. The log-odd ratio maps (third row) look quite similar at first sight to the difference maps, except in the center where the differences are more manifest. Of particular interest is the comparison of the first and second column, i.e. Cartesian vs. “raw” (first column), and Retinotopic vs. “raw” (second column). The radius of the central spot appears clearly different in the two cases, with a large central spot in the Cartesian vs. raw case, reflecting a loosy spatial discrimination improvement, and a more reduced one in the retinotopic vs. raw case, reflecting a more local and object-centric spatial discrimination improvement. Moreover, only the retinotopic network (second column) has significant peripheral depletion in comparison with the “raw” network. These observations collectively support the idea that the retinotopic network generates a heat map with enhanced contrast around the area of interest. In addition though, a focal Cartesian network also manifests good object localization capabilities, with high likelihood values around the object of interest, though being less spatially specific.
Comparing likelihood maps and Intersection over Union
@@
To further quantify the difference between the Cartesian and retinotopic cases, we considered the localization information provided by the bounding boxes. A bounding box is defined as the rectangle of minimal surface containg one whole visual object. As such, each bounding box partitions the image into two regions, a region where the object of interest is present and a region where it is absent. Thus, given each likelihood map in the validation set, we compute the mean likelihood for the label under study within the bounding boxes (ground truth from the dataset) and the mean likelihood outside these boxes. Given the higher classification accuracies of the networks when fine-tuned on bounding boxes only the networks retrained on the “focus” dataset are used for the remainder of the study. The results are shown in Figure 7-A. The figure shows a mean likelihood value obtained inside and outside the bounding boxes for the original ResNet network, our re-trained Cartesian network, and the re-trained retinotopic network. Both networks show significantly higher likelihoods when the fixation point is inside the bounding box than when it is outside the bounding box. This reflects a higher confidence in the label response. At first glance, the average likelihoods in the three conditions seem quite comparable, although the likelihood values appear slightly higher in the re-trained Cartesian case.
@@
Figure 7: We tested the visual search protocol on the ImageNet validation dataset ($50,000$ images). (A) The mean likelihood across the point of fixation inside the bounding box (“In”) or the point of fixation outside the bounding box (“Out”) and the corresponding ratio of activation. (B) The intersection over union as a function of a threshold applied on the likelihood map.
@@
In addition, we consider in Figure 7-B the Intersection over Union (IoU) metric to evaluate the agreement between the bounding boxes and the activation maps (see Methods “Localization tools and evaluation”). In contrast with the “raw” network, our two networks fine-tuned on the “focus” datasets show a slower decay rate, reflecting a better fit with the ground truth. As expected from previous remarks, the IoU stays constistently higher for the cartesian network than for the retinotopic one, illustrating a tendancy for the retinotopic networks to “concentrate” likelihoods on smaller portions of the bounding box. The peak IoU is obtained for likelihood thresholds close to 0 for all networks, this could be an artifact of the measurement, since we are using the bounding box as a comparison, which encourages a larger number of positions with increased likelihood values on the map, rather than a sharp contour around the object of interest.